Two astronauts are floating motionless in space when they push off from each other. Astronaut A has a large mass M, and Astronaut B has a small mass m. After they finish pushing Astronaut B is moving with speed v. Which Astronaut has a larger kinetic energy after the push? (a) Astronaut A (b) Astronaut B (c) They have the same kinetic energy (d) It’s not possible to tell without knowing the numbers
Answer #1

from the conservation of momentum

momentum of A = momentum of B

M*v = M*v'

v' = Mv/m

astronaut A has kinetic energy = kA = 1/2*M*v^2

astronaut B has kB = 1/2*m(Mv/m)^2

kB = 1/2*m(M^2*v^2/m^2)

kB = 1/2*(M/m)*v^2

when compared to both kinetic energies astronaut has more kinetic energy